OpenStudy (anonymous):

If f(x) = 2x2, and 2f(3a) = 144, then what could be the value of a ?

jimthompson5910 (jim_thompson5910):

\[\large f(x) = 2x^2\] \[\large 2*f(x) = 2*2x^2\] \[\large 2*f(x) = 4x^2\] \[\large 2*f(3a) = 4(3a)^2\] \[\large 144 = 4(3a)^2\] I'll let you finish up.